Circular No. 7118

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE Further to IAUC 7117, A. Kim et al. add: “Goobar and Hook obtained

a CCD spectrum of SN 1999ah on Feb. 13 at the NOT; they, along with Nugent and Aldering, find that the host galaxy has z = 0.080 from O II and Hα lines. The spectrum, while heavily contaminated by the underlying host galaxy, displays broad supernova-like emission features at 542 and 592 nm, and is inconsistent with a type-Ia or type-II identification at the redshift of the host galaxy and bears no resemblence to the type-Ic spectra of SNe 1994I or 1983V. SN 1999ai was confirmed with a CCD spectrum taken at the APO 3.5-m telescope by Fan and Newberg on Feb. 17; Nugent, Aldering, and Deustua note that the spectrum is that of a type-II supernova, a couple weeks past peak brightness, at a redshift of 5340 km/s (consistent with that of the host galaxy at z = 0.018) — in agreement with a CCD spectrum taken by Goobar on Feb. 21 at the NOT and analyzed by Hook. Goobar took a CCD spectrum of SN 1999ak on Feb. 21 at the NOT, and he (along with Nugent, Aldering, and Hook) notes that the spectrum of the supernova and host galaxy place both at a redshift of z = 0.055; the supernova spectrum is consistent with that of a type-Ia event, 2 weeks past peak brightness. SN 1999al was confirmed with a CCD spectrum taken at the Lick 3.0m telescope by Aldering on Feb. 25; Nugent and Deustua find that the spectrum is consistent with a type-Ic supernova at z = 0.065, and it is very similar to the earliest spectrum of SN 1983V when it was roughly 9 days before peak brightness (Clocchiatti et al. 1997, Ap.J. 483, 675).”

U SCORPII T. Zwitter, Department of Physics, University of Ljubljana; and U.

Munari, Padova and Asiago Astronomical Observatories, write: “Highresolution spectra (range 430–875 nm) of the recurrent nova U Sco (cf. IAUC 7113) were obtained on Feb. 27.2, 28.2, Mar. 1.2, and 2.2 UT with the Asiago 1.82-m telescope (+ CCD échelle spectrograph). The spectra are remarkably similar over the four nights. Only subtle profile changes are seen in N II, N III, C III, C IV, and He I emission lines. On the other hand, Balmer lines changed appreciably from saddle-like to single-peaked profiles. The FWZI of Hα has remained constant at 8900 km/s. Narrow emission components have not yet appeared.”

Visual magnitude estimates: Feb. 25.04 UT, [14.3 (B. Monard, Pretoria, S. Africa); 26.162, 8.7 (P. Schmeer, Bischmisheim, Germany); 26.794, 9.2 (A. Pearce, Nedlands, W. Australia); 27.144, 9.5 (Schmeer); 28.190, 10.1 (Schmeer); Mar. 1.840, 10.5 (Pearce); 2.873, 11.2 (Pearce).

1999 March 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green